### v1

@@ 4.1 Framework Overview
Based on existing studies[7] and following discussions with multiple physicians, we further decomposed the first three tasks in the pre-consultation process into several domain-specific subtasks (see Table 5). Unlike the decomposable collection tasks, CC generation ($T_{4}$) is a comprehensive integrative task that is not suitable for further subdivision. It operates directly on the Recipient’s incremental recording process, refining information through the continuous accumulation of dialogue content and the dynamic evolution of the HPI. Ultimately, it produces a CC consistent with the full clinical narrative.
Building on the hierarchical task modeling described above, the introduction of a multi-agent framework provides a rational implementation approach. Tasks in the pre-consultation scenario involve both decomposable, structured collection components and context-dependent, integrative generation components, which are often difficult to address simultaneously with a single model. Through the collaborative mechanism of multiple agents, different tasks can be handled by independent autonomous entities while achieving global coordination during interactions, thereby better balancing efficiency and flexibility. The overall architecture is illustrated in Figure 1.
(a) GPT-OSS
(b) Qwen3
(c) Phi4
(d) Uncompleted cases
At the implementation level, we designed eight core agents covering global scheduling, question–answer generation, task evaluation, and other processes. This configuration strikes a balance between task coverage and system complexity: the number of agents is sufficient to allow fine-grained task handling while avoiding redundancy and information fragmentation caused by excessive decomposition. Each agent has a clearly defined functional role, forming a structure of role specialization (see Table 6 for the overall design).
To further illustrate the interactions among the eight agents during task execution, we present an overall system architecture diagram (see Figure 1). The diagram is based on the hierarchical task structure and depicts the execution pathways of $T_{1}$, $T_{2}$, $T_{3}$, and $T_{4}$ under multi-agent collaboration. It also visually represents the information flow between global scheduling and local collection processes.
4.2 Task Scheduling and Monitoring
When the system receives the $k$-th round of updated patient information, which includes Triage, History of Present Illness (HPI), Past History (PH), and Chief Complaint (CC), the current task group for evaluation is selected based on the completion order of different sections. If the current task group remains consistent with the previous round’s task group, the system inherits the pending task set $T_{P}$ from the previous round. The Monitor then re-evaluates each subtask within the pending task set $T_{P}$ by incorporating the newly acquired patient information to generate updated Triage, HPI, PH, and CC information. The evaluation is based on two dimensions: clinical semantic validity and task completeness, with scores ranging from 0 to 1. Subtasks that achieve a score above the threshold of 0.85 are removed from the pending task set $T_{P}$, while those scoring below this threshold remain in the set for further processing.
$T_{P}^{k+1}=\{t_{ij}|S(t_{ij})<0.85,t_{ij}\in T_{P}^{k}\}$ (5)
where $S(\cdot)$ denotes the evaluation function of the Monitor, which assesses each subtask based on clinical semantic validity and task completeness.
In this way, the role of the Monitor is not only to evaluate task quality but also to balance information integrity and dialogue efficiency through the threshold mechanism: an excessively high threshold may lead to redundant follow-up questions and over-collection of information, while a threshold that is too low risks missing critical clinical details. Once the Monitor completes its evaluation, the Controller $C$ performs global scheduling based on the real-time state of the pending tasks. It integrates task priority, dialogue context, and the patient’s current clinical information continuously updated by the Recipient. From the pending task set, it selects the subtask with the highest priority as the current subtask $T_{C}$ and generates corresponding inquiry instructions. These instructions are then passed to the Prompter and Inquirer, thereby triggering targeted follow-up questioning.
$T_{C}^{k}=C(T_{p})^{k}$ (6)
4.3 Information Update and Question Generation
Task scheduling alone is insufficient to support a fully interactive consultation. While the scheduling mechanism specifies the current task $T_{C}$, the system must dynamically update patient information during dialogue and generate targeted clinical questions. Without this closed-loop, even accurately assigned tasks cannot ensure the professionality or coherence of the inquiry. To address this, we introduce an information updating and question generation framework atop task scheduling.
Upon receiving natural language expressions from the Virtual Patient, the Recipient integrates the new information into the existing medical record and incrementally updates the Chief Complaint (CC), History of Present Illness (HPI), and Past History (PH). This mechanism allows patient history to evolve across dialogue turns rather than remain at the level of isolated statements, providing a stable context for subsequent tasks. Formally:
$(\mathrm{HPI}_{k+1},\mathrm{PH}_{k+1},\mathrm{CC}_{k+1})=F(\mathrm{HPI}_{k},\mathrm{PH}_{k},\mathrm{CC}_{k},\mathcal{C}_{k+1}),$ (7)
where $\mathcal{C}_{k+1}$ is the complete dialogue up to turn $k+1$, and $F$ denotes the Recipient’s update function. Iterative application ensures progressive accumulation of medical record information.
After updating HPI, PH, and CC, the Prompter generates guiding prompts for the Inquirer to enable more targeted questioning. The Prompter integrates multiple inputs: the updated HPI, PH, and CC from the Recipient, and importantly, the current task $T_{C}$ provided by the Controller. By incorporating $T_{C}$, the Prompter ensures that the generated prompts align precisely with the current task objectives while remaining within the scope of clinician–patient interactions, avoiding irrelevant or redundant queries. In effect, the Prompter operationalizes the abstract goals of task scheduling.
Guided by the prompts from the Prompter and the updated medical record, the Inquirer produces context-sensitive clinical questions. For a given current task $T_{C}$, the Inquirer’s outputs reflect the scheduling objectives while incorporating the constructed clinical context, maintaining both logical continuity and medical professionalism.
Additionally, when task $T_{1}$ has not been completed, the Triager is invoked as an auxiliary component to analyze symptom characteristics and provide triage recommendations. It constructs a contextualized clinical interpretation based on structured taxonomies and outputs candidate recommendations for primary and secondary departments. This maps unstructured dialogue into a clinically interpretable feature space during the initial triage phase. Once $T_{1}$ is completed, the Triager is no longer called in subsequent rounds.
4.4 Evaluation
Within the entire multi-agent framework, the Evaluator is designed as an independent assessment agent responsible for systematically measuring system performance. It is important to emphasize that the Evaluator does not participate in the actual clinician–patient dialogue and does not influence task scheduling or subtask selection; its function is entirely limited to post hoc evaluation. Thus, it serves a role of performance characterization and quality feedback rather than interactive execution. To ensure comprehensive and rigorous evaluation, the Evaluator operates on two complementary levels: dialogue interaction and content accuracy. This distinction reflects the dual objectives of clinical pre-consultation: ensuring both the rationality of the interaction process and the accuracy of the generated content.
At the interaction level, it examines whether the system’s questions follow clinical logic and task objectives, the clarity and professionalism of language, the completeness of information collection, and the overall professionalism demonstrated during the inquiry. At the content level, it evaluates the alignment of the three key medical record elements—Chief Complaint, History of Present Illness, and Past History—against reference standards to ensure consistency and fidelity. Detailed definitions and criteria for each dimension are provided in Table 7.
Regarding scoring methodology, the Evaluator uses a 0–5 scale for all seven dimensions, where 0 represents severe omission or error, and 5 represents an ideal clinical level. Complementarily, qualitative criteria are provided to interpret each score range with specific explanations and examples. For instance, Table 8 presents the scoring standards for Past History Similarity (PHS), considering information coverage, chronological consistency, and accuracy of medical terminology. By combining quantitative and qualitative assessments, the Evaluator not only ensures objective scoring but also provides interpretable feedback for subsequent system optimization.
5 Experiment
5.1 Triage Experiment
To evaluate the effectiveness of our multi-agent framework in medical triage scenarios, we conducted comprehensive experiments on department classification tasks. We assessed the system’s ability to correctly route patients to both primary and secondary departments based on their presenting symptoms and medical history.
Table 2 presents the classification accuracy across multiple iteration steps. Our results demonstrate consistent improvement in triage accuracy through iterative refinement. The primary department classification accuracy increased from 83.0% in the initial step to 87.0% by step 4, representing a 4.0 percentage point improvement. Similarly, secondary department classification showed substantial gains, rising from 75.4% to 80.5%, an improvement of 5.1 percentage points. The progressive improvement across iterations validates our multi-agent approach, where agents collaboratively refine their decisions through structured dialogue and knowledge sharing.
Table 2: Performance progression across steps for primary and secondary department classification. Results show accuracy improvements through multi-round iteration.
Iteration Step
Primary
Department
Secondary
Department
Step 1 83.0% 75.4%
Step 2 85.6% 78.7%
Step 3 86.4% 80.2%
Step 4 87.0% 80.5%
Figure 3 illustrates the triage accuracy distribution across different medical departments. We observe significant variability in performance across specialties, with accuracy ranging from approximately 65% to 95%. Ophthalmology achieved the highest classification accuracy at 94.8%, likely due to the distinctive nature of ocular symptoms that facilitate clear differentiation. In contrast, Psychiatry exhibited the lowest accuracy at 65.2%. Further analysis revealed that misclassified psychiatry cases were predominantly routed to the neurology department. This pattern likely reflects the inherent overlap between neurological and psychiatric presentations, as well as potential variations in departmental structures across different hospital systems.
Figure 3: Triage accuracy across different medical departments. The bar chart shows significant variation in classification performance, with Ophthalmology achieving the highest accuracy (100.0%) and Psychiatry the lowest (30.0%).
These findings demonstrate the value of our multi-agent architecture in handling the inherent ambiguity of medical triage. The consistent improvement through iterative refinement and the ability to achieve high accuracy for departments with distinct symptom profiles suggest our system could provide immediate value in clinical settings, particularly in routing clear-cut cases and flagging ambiguous cases for human review.
5.2 Experimental Design
All experiments in this study were conducted based on our self-constructed dataset comprising 1,372 validated medical records. These records serve as the foundation for evaluating the performance of our proposed consultation system across various clinical scenarios. For model deployment, all models were deployed on a single NVIDIA A100 GPU using GPUStack as the deployment framework. To enhance inference efficiency, we implemented concurrent processing with 60 parallel threads, allowing the system to handle multiple consultation sessions simultaneously while maintaining response quality. The maximum number of dialogue turns was set to 30 rounds per consultation session. This constraint ensures practical applicability while providing sufficient opportunity for comprehensive information gathering. Samples that failed to complete all three task groups ($T_{1}$, $T_{2}$, and $T_{3}$) within the 30-round limit were classified as failures in our evaluation metrics. To ensure consistency and fairness in evaluation across all experiments, we standardized the Evaluator component by uniformly employing the GPT-OSS 20B model. This standardization eliminates potential variations that could arise from using different evaluation models, thereby ensuring reliable and comparable results throughout our experimental analysis.
5.3 Task completion across different models
To evaluate the generalizability of the hierarchical multi-agent architecture, we conducted zero-shot tests on three LLMs of different scales: GPT-OSS 20B, Qwen3-8B, and Phi4-14B. The results (Figure 2) show that our framework can guide all models to complete multi-turn medical consultations without any task-specific fine-tuning, highlighting its model-agnostic nature.
Specifically, the powerful GPT-OSS 20B achieved the most stable and reliable performance under the framework’s coordination, with the lowest number of unfinished cases. The combined unfinished count for $T_{2}$ and $T_{3}$ tasks by GPT-OSS 20B was only 55.43% of that achieved by Qwen3-8B and merely 8.08% of Phi4-14B’s unfinished cases, demonstrating superior task completion capability. Interestingly, Qwen3-8B demonstrated remarkable efficiency in task execution. It completed both $T_{2}$ and $T_{3}$ tasks with fewer average rounds compared to GPT-OSS 20B. This efficiency advantage may be attributed to Qwen3-8B’s stronger Chinese language capabilities, which are particularly relevant for our medical consultation scenarios. However, its lower completion rate compared to GPT-OSS 20B suggests that superior reasoning abilities remain crucial for ensuring task success. For Phi4-14B, the overall performance lagged behind both GPT-OSS 20B and Qwen3-8B, which may be related to its earlier release date (Phi4-14B was released in 2024, while GPT-OSS 20B and Qwen3-8B were released in 2025). Despite this limitation, the framework still exhibited strong robustness and fault tolerance, successfully guiding Phi4-14B to complete a considerable number of cases rather than failing entirely.
All models required more rounds to complete the more complex $T_{3}$ tasks, which is consistent with expectations and further validates the soundness of the evaluation design. The experimental results demonstrate that the system maintains effectiveness across different models, and notably, task completion success rate does not solely depend on model parameter size but rather correlates more strongly with language-specific capabilities and general reasoning abilities.
5.4 Evaluation scores across different models
As illustrated, all three models achieved commendable evaluation scores across the six distinct medical consultation tasks (CI, CQ, IC, OP, PHS, CCS). The consistent presence of high median values (black line within each box) and the tight interquartile ranges (compact box size) indicate that our architecture effectively guides different LLMs to perform the structured reasoning required for multi-round dialogues.
In the three metrics that reflect information extraction effectiveness—Chief Complaint Similarity (CCS), Present Illness Similarity (HPIS), and Past History Similarity (PHS)—all three models (GPT-OSS 20B, Qwen3-8B, and Phi4-14B) achieved average scores of at least 3.74, as shown in Figure 4. Specifically, GPT-OSS 20B achieved an average of 3.99, Qwen3-8B scored 3.74, and Phi4-14B also reached 3.99. These consistently high scores demonstrate that our system can achieve excellent final outcomes across different models, regardless of their individual characteristics.
Regarding the four metrics that reflect dialogue process quality (CI, CQ, IC, and OP), Phi4-14B demonstrated slightly stronger performance compared to GPT-OSS 20B and Qwen3-8B. This performance variation may be attributed to differences in the linguistic styles of different models, suggesting that certain conversational characteristics inherent to Phi4-14B align well with the dialogue evaluation criteria.
5.5 Completion rate across different scheduling strategies
We compare the dynamic performance of two representative scheduling strategies: Default Order (baseline sequential prompting following a predefined sequence, which we designed with reference to clinical importance) and Agent Driven (autonomous coordination through agent selection).
As shown in Fig. 5, both strategies exhibit a monotonic upward trend in task completion across the 13 subtasks. However, they differ in convergence levels and growth rates.
Specifically, the Agent Driven strategy demonstrates the fastest growth in the early phase and ultimately stabilizes at the highest completion rate of 98.2% around round 25, reflecting superior global coordination and convergence capacity. In contrast, the Default Order strategy maintains a steadier growth trajectory during the middle phase (rounds 5–15) but plateaus at a lower final completion rate of 93.1%, suggesting a performance ceiling.
These results indicate that the Agent Driven strategy achieves a more favorable balance between stability and efficiency, enabling faster and more complete subtask fulfillment. Meanwhile, the Default Order approach provides moderate consistency but lacks the capacity to overcome performance saturation. The superior performance of Agent Driven strategy demonstrates the advantage of dynamic, intelligent scheduling over static sequential processing, even when the default sequence is designed with clinical considerations in mind.
(a) Agent Driven
(b) Medical Priority
Figure 5: Task completion rates under different scheduling strategies. Our proposed Agent Driven strategy demonstrates superior completion performance compared to the baseline Medical Priority approach.
5.6 Evaluation scores across different scheduling strategies
To further assess the impact of scheduling strategies on medical history generation, we compared Default Order and Agent Driven across six clinically relevant dimensions: Chief Complaint Similarity (CCS), Present Illness Similarity (HPIS), Past History Similarity (PHS), Inquiry Consistency (IC), Overall Progress (OP), and Clinical Integrity (CI).
As illustrated in Fig. 6, the Agent Driven strategy consistently outperforms Default Order across most dimensions. It achieves notable gains in IC and OP, reflecting stronger coherence and broader coverage, and shows clear advantages in PHS and HPIS, suggesting more complete and accurate medical history collection. While the two strategies achieve comparable performance in CCS, Agent Driven still maintains a slight edge. The only area where Default Order remains competitive is CI, though its overall balance lags behind.
Taken together, these results highlight the robustness of Agent Driven scheduling: it not only accelerates task completion (as shown in the previous subsection) but also produces more consistent and comprehensive outputs across diverse evaluation metrics. This demonstrates that dynamic, intelligent coordination surpasses static sequential processing, even when the default sequence is carefully designed based on clinical considerations.
Figure 6: Evaluation scores across different scheduling strategies
5.7 Real-World Evaluation
Although the Evaluator-based scoring offers consistency and reference value, its evaluation focus may differ from real-world pre-consultation settings. To validate the usability and clinical relevance of the results generated by the Recipient, we incorporate expert physician evaluation as a complementary metric. A total of $40$ samples were randomly selected from a dataset of $1,372$ records. For each sample, considering that lengthy dialogues may cause evaluation fatigue and affect scoring consistency, we randomly sampled $5$ out of the $20$ dialogue turns to present to evaluators. This sampling strategy balances the need for comprehensive evaluation while maintaining physician focus and reducing cognitive load. The sampled dialogue turns, along with the final-round generated Chief Complaint (CC), History of Present Illness (HPI), and Past History (PH) were presented. A total of $18$ licensed physicians were invited to participate in the evaluation, with each physician independently assessing $20$ samples. Each sample was thus evaluated by $9$ physicians, who assigned a score on a discrete scale from $1$ to $5$.
Table 3: Real World Evaluation
Field Average Score Standard Deviation
$C$ 4.5647 0.2221
HPI 4.4794 0.1284
PH 4.6853 0.0997
Table 3 presents the expert physician evaluations on generated content across different fields. The average scores are above $4.45$, indicating generally high quality, while relatively low standard deviations (below $0.23$) demonstrate consistent agreement among raters.
Acknowledgments
Appendix A Case Study and Human Evaluation
Next, we present a representative case study featuring a 5-round dialogue transcript along with the corresponding generated HPI, PH, and CC outputs. Additionally, we provide evaluation scores from nine medical practitioners for this case to better demonstrate the practical effectiveness of our proposed model.
A.1 Dialogue Excerpt
Patient-Doctor Consultation Transcript Patient: My neck and shoulder have been hurting for half a month. It’s hard to move. Doctor: Please think back carefully. When the pain started, did anything specific happen? Patient: That night, I got up too quickly while turning over, and suddenly my right neck cramped with sharp pain. After resting for 10 minutes, it eased. Doctor: Can you describe the nature of the pain? Is it a dull constant ache, intermittent stabbing, or something else? Patient: It’s mostly spasmodic pain in the neck and right shoulder. Very intense during attacks, but subsides after rest. The left shoulder feels sore and stiff in the morning. Doctor: Any other symptoms, such as numbness in the arm, dizziness, or vision changes? Patient: No, just the neck and shoulder pain. Doctor: Have you ever been diagnosed with any chronic diseases, like hypertension, diabetes, heart disease, hepatitis, etc.? Patient: No, I’ve always been healthy. Doctor: Any allergies to drugs or foods? Patient: No known allergies.
A.2 Generated Documentation
System-Generated Medical Documentation Chief Complaint ($C$):
Neck and shoulder pain with limited mobility for half a month.
History of Present Illness (HPI):
Patient developed neck and shoulder pain with restricted mobility half a month ago. One day ago, a sudden movement while getting up led to acute cramping pain in the right neck, which resolved after 10 minutes of rest. The pain is intense during attacks and alleviates with rest. Left shoulder shows noticeable morning soreness and limited mobility. Denies numbness or dizziness.
Past History (PH):
Denies any history of chronic diseases such as hypertension or diabetes. No prior surgeries. No known drug or food allergies.
A.3 Human Evaluation Results
Nine medical practitioners evaluated the generated documentation quality, with scores shown in Table 4.
Table 4: Expert evaluation scores on one representative case
Field 1 pt 2 pt 3 pt 4 pt 5 pt Avg. Score
$C$ 0 0 0 2 7 4.78
HPI 0 0 0 3 6 4.67
PH 0 0 0 0 9 5.00
Appendix B Supplementary Tables and Figures
This appendix provides detailed task definitions and additional visualizations that complement the main text. The hierarchical task decomposition framework (Table 5) presents the complete structure of our pre-consultation system, while supplementary figures and tables offer comprehensive evaluation criteria and data distributions.
B.1 Task Decomposition Framework
The following table presents the complete hierarchical decomposition of pre-consultation tasks, which forms the foundation of our multi-agent coordination system.
@@ B.1 Task Decomposition Framework
B.2 Data Distribution and System Architecture
Figure 7 illustrates the distribution of medical cases across primary departments in our validation dataset, demonstrating the comprehensive coverage of clinical specialties.
B.3 Agent Architecture and Evaluation Framework
The following tables provide detailed specifications of our multi-agent system components and evaluation criteria used throughout the experimental validation.

### v2

@@ 2.2.1 Task Formalization
where $T_{1}$ represents Triage, $T_{2}$ represents HPI collection, $T_{3}$ represents PH collection, and $T_{4}$ represents CC generation. This hierarchical modeling approach facilitates a balance between global optimality and efficient completion of individual tasks.
@@
Based on existing clinical guidelines [gawlik2024historytaking] and discussions with multiple physicians, we decomposed the first three tasks into domain-specific subtasks (Table 1). Unlike the decomposable collection tasks, CC generation ($T_{4}$) is a comprehensive integrative task that operates directly on the incremental recording process, refining information through continuous accumulation of dialogue content and dynamic evolution of the HPI.
2.2.2 Multi-Agent Architecture
Figure 2: Hierarchical multi-agent framework architecture for medical consultation workflow. Parts 1-3 correspond to Sec. 2.2.3, 2.2.4 respectively.
The multi-agent framework addresses the challenge that pre-consultation tasks involve both decomposable, structured collection components and context-dependent, integrative generation components, which are difficult to handle simultaneously with a single model. Through collaborative mechanisms, different tasks are handled by independent autonomous entities while achieving global coordination during interactions. The overall architecture is illustrated in Figure 2.
@@ 2.2.2 Multi-Agent Architecture
We designed eight core agents covering global scheduling, question–answer generation, and task evaluation. This configuration balances task coverage and system complexity: the number of agents is sufficient to allow fine-grained task handling while avoiding redundancy and information fragmentation. Each agent has a clearly defined functional role (Table 2).
2.2.3 Task Scheduling and Monitoring
When the system receives the $k$-th round of updated patient information (including Triage, HPI, PH, and CC), the current task group for evaluation is selected based on the completion order of different sections. If the current task group remains consistent with the previous round, the system inherits the pending task set $T_{P}$ from the previous round.
The Monitor agent evaluates each subtask within the pending task set $T_{P}$ by incorporating newly acquired patient information. Evaluation is based on two dimensions: clinical semantic validity and task completeness, with scores ranging from 0 to 1. Subtasks achieving scores above the threshold of 0.85 are removed from the pending task set, while those scoring below remain for further processing:
$T_{P}^{k+1}=\{t_{ij}\mid S(t_{ij})<0.85,t_{ij}\in T_{P}^{k}\}$ (5)
where $S(\cdot)$ denotes the evaluation function of the Monitor. The threshold mechanism balances information integrity and dialogue efficiency: an excessively high threshold may lead to redundant follow-up questions, while a threshold that is too low risks missing critical clinical details.
Once the Monitor completes evaluation, the Controller agent performs global scheduling based on real-time pending task states. It integrates task priority, dialogue context, and patient clinical information to select the highest-priority subtask as the current task $T_{C}$:
$T_{C}^{k}=C(T_{P}^{k})$ (6)
The Controller then generates corresponding inquiry instructions passed to the Prompter and Inquirer agents.
2.2.4 Information Update and Question Generation
Task scheduling alone is insufficient to support fully interactive consultation. The system must dynamically update patient information during dialogue and generate targeted clinical questions to ensure professionality and coherence of inquiry.
Upon receiving natural language expressions from the patient, the Recipient agent integrates new information into the existing medical record and incrementally updates CC, HPI, and PH:
$(\mathrm{HPI}_{k+1},\mathrm{PH}_{k+1},\mathrm{CC}_{k+1})=F(\mathrm{HPI}_{k},\mathrm{PH}_{k},\mathrm{CC}_{k},\mathcal{C}_{k+1})$ (7)
where $\mathcal{C}_{k+1}$ is the complete dialogue up to turn $k+1$, and $F$ denotes the Recipient’s update function. This mechanism allows patient history to evolve across dialogue turns rather than remain at the level of isolated statements.
The Prompter agent generates guiding prompts for targeted questioning by integrating updated HPI, PH, and CC from the Recipient with the current task $T_{C}$ from the Controller. This ensures generated prompts align precisely with current task objectives while remaining within the scope of clinician–patient interactions.
Guided by prompts from the Prompter and the updated medical record, the Inquirer agent produces context-sensitive clinical questions reflecting scheduling objectives while maintaining logical continuity and medical professionalism.
Additionally, when task $T_{1}$ (Triage) has not been completed, the Triager agent analyzes symptom characteristics and provides triage recommendations by constructing contextualized clinical interpretations based on structured taxonomies. Once $T_{1}$ is completed, the Triager is no longer invoked.
2.2.5 Safety Constraints in Agent Design
To mitigate potential risks associated with LLM-generated content in medical contexts, we implemented prompt-level safety constraints across all agents. Specifically, the Recipient and Inquirer agents were explicitly instructed to (1) refrain from fabricating or inferring medical information not directly stated by the patient, (2) avoid leading or suggestive questioning that could bias patient responses, and (3) acknowledge uncertainty when information is incomplete rather than generating plausible but unverified content. These constraints were reinforced through few-shot examples demonstrating both appropriate and inappropriate agent behaviors within the system prompts. However, we acknowledge that this approach relies on prompt-based guidance rather than architectural enforcement mechanisms.
2.3 Framework Evaluation Methods
The performance of the proposed framework was assessed using both quantitative metrics and qualitative evaluation dimensions.
2.3.1 Quantitative Assessment: Interaction and Content Quality
Within the multi-agent framework, the Evaluator is designed as an independent assessment agent responsible for systematically measuring system performance. The Evaluator does not participate in actual clinician–patient dialogue and does not influence task scheduling or subtask selection; its function is entirely limited to post hoc evaluation.
To ensure comprehensive and rigorous evaluation, the Evaluator operates on two complementary levels reflecting the dual objectives of clinical pre-consultation:
Interaction-Level Assessment
At the interaction level, the Evaluator examines: (1) whether system questions follow clinical logic and task objectives, (2) clarity and professionalism of language, (3) completeness of information collection, and (4) overall professionalism demonstrated during inquiry.
Content-Level Assessment
At the content level, the Evaluator assesses alignment of the three key medical record elements—Chief Complaint, History of Present Illness, and Past History—against reference standards to ensure consistency and fidelity.